OpenStudy (anonymous):

Multiply 2+3i/-2i 1. - 15/19- 15/38 i 2. -22/9+ 71/18 i 3. 135/92- 3/46 i 4. -3/2 + i 5. None of the above

OpenStudy (anonymous):

oh nvm I got it, it's number 4

OpenStudy (xapproachesinfinity):

you mean simplify? by rationalizing the denominator

OpenStudy (xapproachesinfinity):

actually it's 4

OpenStudy (xapproachesinfinity):

it is not 4 i mean

OpenStudy (xapproachesinfinity):

\(\large \frac{2+3i}{-2i}=-3/2-i\)

OpenStudy (anonymous):

oh I see, thank you

OpenStudy (xapproachesinfinity):

YW!